a horizontal Force 0.80N is required to pull a 5 kg block across a table top at constant speed with the block to slide home 1.5m before coming to rest again .determine the speed v of the bullet ,where the bullet is assumed to be embedded in a block
1
Expert's answer
2021-08-16T08:36:54-0400

Find the initial speed of the block before it started slowing down along a 1.5m-distance:


"v=\sqrt{2ad}=\sqrt{2\frac FMd}=0.48\text{ m/s}."

The speed of the bullet can be found by momentum conservation assuming its mass is infinitesimal compared to the block:


"mu=Mv,\\\space\\
u=\frac Mmv=\frac{2.4}{m}."